OpenStudy (anonymous):

he output voltage of a transistor switching circuit is described by the following Fourier series: v=(2/π)∑[[((-1)^n) /n][cos(nπ/2)-1]sin(nπt/2)] (for n=1 to∞) 1. What is the amplitude of the second harmonic of this voltage? 2. What is the DC component of this voltage? 3. What is the fundamental frequency of the output

OpenStudy (anonymous):

\[v=\left[ \frac{2}{\pi} \right]\Sigma \left[ \frac{-1^n}{n} \right]\left[ \cos \left( \frac{n \pi}{2} -1\right) \right]\left[ \sin \frac{n \pi t}{2} \right]_{1}^{\infty}\]Is this how the equation should look?